OpenStudy (anonymous):

please help me everyone . what is the derivative of: f(t)=20[1-cos^2(t/1000pi)]

OpenStudy (anonymous):

Put 1-cos^2(t/1000pi) = sin^2(t/1000pi) and start differentiating using chain rule. Consider 20[sin(t/1000pi)]^2 as the function to start differentiating with.

OpenStudy (abb0t):

\[-\frac{ 20 \pi }{ 1000 }[2\cos (\frac{ t \pi }{ 1000 })\sin(\frac{ t \pi }{ 1000 })]\]